# Discrete Guidance Matching: Exact Guidance for Discrete Flow Matching

### v1

@@ Abstract
Guidance provides a simple and effective framework for posterior sampling by steering the generation process towards the desired distribution. When modeling discrete data, existing approaches mostly focus on guidance with the first-order Taylor approximation to improve the sampling efficiency. However, such an approximation is inappropriate in discrete state spaces since the approximation error could be large. A novel guidance framework for discrete data is proposed to address this problem: We derive the exact transition rate for the desired distribution given a learned discrete flow matching model, leading to guidance that only requires a single forward pass in each sampling step, significantly improving efficiency. This unified novel framework is general enough, encompassing existing guidance methods as special cases, and it can also be seamlessly applied to the masked diffusion model. We demonstrate the effectiveness of our proposed guidance on energy-guided simulations and preference alignment on text-to-image generation and multimodal understanding tasks. The code is available through https://github.com/WanZhengyan/Discrete-Guidance-Matching/tree/main.
@@ 1 Introduction
Discrete diffusion models (austin2021structured, ; campbell2022continuous, ; sun2022score, ; lou2023discrete, ) and discrete flow-based models (campbell2024generative, ; gat2024discrete, ; shaul2024flow, ; yimingdefog, ) have received significant attention for generating samples in discrete state spaces, providing effective alternatives to autoregressive (AR) models. Therefore, several works aim to study how to guide a pre-trained model towards the desired distribution (vignac2022digress, ; schiff2024simple, ; nisonoff2024unlocking, ). However, there are some challenges to developing guidance mechanisms in discrete cases. The main challenge is that discrete guidance is often associated with a transition probability or transition rate, which requires extra computation for all possible target positions after the transition. To reduce the number of forward passes through guidance models, existing methods treat the model as a continuous function and use the first-order Taylor approximation to compute guidance efficiently (vignac2022digress, ; schiff2024simple, ; nisonoff2024unlocking, ). Unfortunately, this approximation might introduce non-negligible approximation errors. Moreover, existing literatures only consider some particular cases, like class conditional guidance or energy-weighted guidance, which might not be general enough for various tasks.
In this article, we seek to develop a general guidance framework for discrete diffusion and flow-based models, which is shown in Fig. 1. Under the Continuous-Time Markov Chain (CTMC) settings, given a pretrained model for sampling from a source distribution and the density ratio between the target distribution and source distribution, our method identifies the exact transition rate for the target distribution, i.e., a conditional expectation of the density ratio between the target and source distribution. To the best of our knowledge, our guidance framework is the first exact discrete guidance in general form. Since our guidance formulation is expressed by the density ratio between the target distribution and source distribution, it encompasses existing guidance methods as special cases. The density ratio can be calculated by the energy function in energy-guided sampling, the ratio of the classifier in class-conditional generation, and preference alignment in reinforcement learning from human feedback (RLHF). In Table 1, we compare the formulation of the guidance and the number of function calls in each sampling step. Our framework only requires one forward pass in each sampling step for any initial distribution.
To calculate the proposed guidance, we employ the Bregman divergence to train a network for estimating the conditional expectation. To utilize the sample from the target distribution, we further propose a regularization term.
@@
    We introduce a guidance framework for discrete flow matching, achieving efficient sampling without approximation.
@@
    We propose to learn the guidance network by minimizing the Bregman divergence. We further propose a regularization technique to utilize the sample from the target distribution.
@@
    We verify the effectiveness of the proposed framework on energy-based sampling through simulations and preference alignment on text-to-image and multimodal understanding benchmark.
Table 1: Overall comparison with existing discrete guidance. Our posterior-based guidance provides an exact formulation of guidance for general target distributions with a sampling-efficient implementation. Existing methods are either limited to class-conditional generation or suffer from non-negligible approximation errors with multiple function calls in each sampling step. A more detailed discussion is provided in Section 3.
Guidance Method Formula Exact Guidance # of Function Calls
@@
Rate-Based nisonoff2024unlocking Equation 9
Absorbing: $\mathcal{D}+1$
Uniform: $\mathcal{D}\times({|\mathcal{S}|}-1)+1$
First-Order Approximated nisonoff2024unlocking
vignac2022digress ; schiff2024simple Equation 13 2
@@ 2 Discrete Flow Models
Discrete flow models (DFMs) aim to generate samples from a $\mathcal{D}$-dimensional discrete state-space with probability mass function (pmf) $q_{1}$, defined on the domain $\mathcal{S}^{\mathcal{D}}$, where $\mathcal{S}$ denotes the finite state space of each dimension. A natural example of this setting is text generation, where each coordinate corresponds to a token drawn from the vocabulary $\mathcal{S}$. To begin with, we first introduce some background on discrete flow models (campbell2024generative, ; gat2024discrete, ; shaul2024flow, ) under the framework of CTMC (norris1998markov, ), which we describe in detail in Appendix C.
@@ 2.1 Transition Rate and Transition Probability
Thus, the transition rate $u_{t}^{q}$ is also the generator of the CTMC ${\{q_{t}\}}_{t\in[0,1]}$ (holderrieth2024generator, ).
@@ 2.2 Modeling Transition Rate via Marginalization Trick
Given the transition rate $u_{t}^{q}(z,x)$ and current state $x$ at time $t$, we can generate a sample from $q_{t+h}$ using the equation 2, with a sufficiently small time step $h$. Therefore, we can learn a transition rate $u_{t}^{q}$ that can transport from an initial noise distribution $q_{0}$ to a target data distribution $q_{1}$. To obtain such a transition rate for sampling, a natural method is to learn the conditional expectation of the conditional transition rate $u^{q}_{t}(z,x|{\mathbf{x}}_{1})$ that generates the conditional probability path $q_{t|1}(\cdot|{\mathbf{x}}_{1})$ interpolating from noise to the datapoint ${\mathbf{x}}_{1}$. Then $u^{q}_{t}(z,x)=\mathbb{E}_{q_{1|t}({\mathbf{x}}_{1}|x)}{[u^{q}_{t}(z,x|{\mathbf{x}}_{1})]}$ can generate the target probability path $q_{t}$ (see Proposition 3.1 of campbell2024generative, ), i.e., it satisfies the Kolmogorov forward equation 1. For completeness, we include the proof in Appendix D.1. We are free to define the conditional probability path $q_{t|1}(\cdot|{\mathbf{x}}_{1})$ and the conditional transition rate as needed.
It is worth noting that learning a ${|\mathcal{S}|}^{\mathcal{D}}$-dimensional vector-valued function ${(u_{t}(z,x))}_{z\in\mathcal{\mathcal{S}}^{\mathcal{D}}}$ of current time $t$ and state $x$ is intractable when $\mathcal{D}$ is relatively large. To handle such a high-dimensional setting, a common approach is to construct a coordinate-wise conditional probability path and transition rate, that is,
@@
which means that the elements of the vector ${\mathbf{x}}_{t}$ are independent conditional on ${\mathbf{x}}_{1}$. Here, $u_{t}^{q,d}(z^{d},x^{d}|x_{1}^{d})$ is the conditional transition rate that generates the conditional probability path $q_{t|1}^{d}$. A popular choice of probability path and the associated conditional transition rate used in the previous works (campbell2024generative, ; gat2024discrete, ) is
@@
After defining the conditional path and rate, the unconditional transition rate is given by
@@
where $q_{1|t}^{d}(x_{1}^{d}|x)=\sum_{x^{\backslash d}_{1}}q_{1|t}(x_{1}|x)$ is the marginal conditional pmf. Therefore, we only need to learn a $\mathcal{D}\times{|\mathcal{S}|}$-matrix-valued function $U_{t}^{q}(x)=(u_{t}^{q,d}(s,x))_{d,s}$. Then we can update each token in parallel during sampling.
@@
which is the cross-entropy between the true posterior $q_{1|t}$ and the estimated posterior $q^{\theta}_{1|t}$ (campbell2024generative, ; gat2024discrete, ; wang2025fudoki, ).
Additionally, there are some alternative approaches to train the transition rate, including directly maximizing the ELBO in terms of the unconditional transition rate (shaul2024flow, ) and minimizing a Bregman divergence-type conditional generator matching loss (holderrieth2024generator, ). Parameterization through the posterior gives us a unified framework to formulate the discrete guidance not only for discrete flow-based models, but also for other generative models. We include the detailed discussion in Section E.1. Moreover, details of the sampling implementation for equation 2 in Algorithm 1 are provided in the Appendix A.
@@ 3 Discrete Flow Guidance
In this section, we introduce a novel guidance framework to model discrete data. This guidance can be applied to both discrete flow models and discrete diffusion models (campbell2022continuous, ; sun2022score, ; austin2021structured, ; lou2023discrete, ; shi2024simplified, ; sahoo2024simple, ). More detailed discussion can be found in Appendix E.
@@ 3.1 Problem Formulation
Consider a CTMC with a probability path $p_{t}$ associated with a transition rate $u^{p}_{t}$. Suppose that the conditional transition rate and the conditional probability path are prespecified (see equation 4 for an example), and we are given a pre-trained posterior $p_{1|t}$ such that the unconditional transition rate $u^{p}_{t}(z,x)=\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1|t}({\mathbf{x}}_{1}|x)}{[u_{t}^{p}(z,x|{\mathbf{x}}_{1})]}$ can generate the probability path $p_{t}$. We will call the distribution with the pmf $p_{1}$ the source distribution. Our goal is to generate samples from the target distribution $q_{1}$ by rectifying the posterior or the unconditional transition rate in terms of the probability path $p_{t}$.
@@ 3.2 Exact Discrete Guidance
Here, we only assume that the conditional probability paths are the same. Theorem 1 provides a general guidance framework for both discrete diffusion and flow models, because both of them can be modeled via parameterizing posterior densities. In discrete diffusion models, if we further assume that the transition rates of both forward noising processes are equal, then we can generalize and recover the predictor guidance in nisonoff2024unlocking , which we state in the following.
@@
which recovers Equation 2 in nisonoff2024unlocking .
@@
In general cases, the guidance is time-dependent according to Theorem 1 and Theorem 2. However, in the masked diffusion and flow models with $x_{1}$-independent scheduler, the guidance is time-independent. To see this, note that in these cases the posterior is time-independent (see Theorem 1 of ou2024your and Proposition 6 of gat2024discrete for details). Since the guidance is the integral of the product of the density ratio and the posterior, the time-independent posterior can lead to a time-independent guidance.
@@ 3.3 Training Objective
According to Theorem 1, given the density ratio $r(x)$, we can obtain the exact guidance term by learning the conditional expectation of the density ratio $h_{t}^{d}(z^{d},x)\overset{\triangle}{=}\mathbb{E}_{{\mathbf{x}}^{\backslash d}_{1}\sim p({\mathbf{x}}^{\backslash d}_{1}|{\mathbf{x}}^{d}_{1}=z^{d},{\mathbf{x}}_{t}=x)}{[r({\mathbf{x}}_{1})]}$. We parameterize the guidance $h_{t}$ using a neural network. A natural training objective for learning conditional expectation is the Bregman divergence (Banerjee2005, ), which is defined as
@@
where $F(\cdot)$ is a convex function. The most common choice of $F$ is $F(x)=\|{x}\|_{2}^{2}/2$, and then the Bregman divergence is the $\ell^{2}$-loss. Unfortunately, as mentioned in lou2023discrete , the $\ell^{2}$-loss does not work well for learning the conditional expectation of a density ratio due to the positive nature of the density ratio. In our work, we take $F(x)=\langle x,\log x\rangle$, and the training objective is given by the following.
@@
The loss $\mathcal{L}_{h,p}$ only requires the source distribution data. Sometimes, we would like to utilize the data sampling from the target distribution to achieve better performance. Similar to ouyang2024transfer , we introduce a regularization term that allows us to learn the guidance $h_{t}$ with sampling ${\mathbf{x}}_{1}\sim q_{1}$ during training. Notice that the minimizer of the objective in equation 5 is the underlying posterior $q_{1|t}$. Combining with Theorem 1, the exact guidance $h_{t}$ is the minimizer of the following objective.
@@ 3.4 Sampling
for each $d\in{[\mathcal{D}]}$, where $h_{t}^{d,\theta}(x)=\sum_{s\in\mathcal{S}}h_{t}^{d,\theta}(s,x)p_{1|t}^{d}(s,x)$. Then, we sample ${\mathbf{x}}^{d}_{t+h}$ by first-order approximation (see equation 14 in Appendix A for details). The overall sampling algorithm can be found in Algorithm 2 in Appendix A and the detailed design for the improved rate-based method proposed in Theorem 2 can be found in Appendix F.
@@
The rate-based guidance involves multiple forward passes, which depend on the number of nonzero elements in the transition rate $u_{t}^{p}(z,{\mathbf{x}}_{t}),~z\neq{\mathbf{x}}_{t}$, given the current state ${\mathbf{x}}_{t}$. If the initial distribution is a masked point mass $\delta_{\mathbbm{m}}$, the rate-based guidance entails $\mathcal{D}+1$ function calls in each sampling step, which is still computationally inefficient; see Fig. 2 (d). For predictor guidance, nisonoff2024unlocking proposed to use the first-order approximation to estimate the log-ratio $\log\frac{p(y|{\mathbf{x}}_{t}=z)}{p(y|{\mathbf{x}}_{t}=x)}$ for efficient sampling. To be specific, we give the general form based on our proposed rate-based guidance:

### v2

@@ Abstract
Guidance provides a simple and effective framework for posterior sampling by steering the generation process towards the desired distribution. When modeling discrete data, existing approaches mostly focus on guidance with the first-order approximation to improve the sampling efficiency. However, such an approximation is inappropriate in discrete state spaces since the approximation error could be large. A novel guidance framework for discrete data is proposed to address this problem: we derive the exact transition rate for the desired distribution given a learned discrete flow matching model, leading to guidance that only requires a single forward pass in each sampling step, significantly improving efficiency. This unified novel framework is general enough, encompassing existing guidance methods as special cases, and it can also be seamlessly applied to the masked diffusion model. We demonstrate the effectiveness of our proposed guidance on energy-guided simulations and preference alignment on text-to-image generation and multimodal understanding tasks. The code is available at https://github.com/WanZhengyan/Discrete-Guidance-Matching.
@@ 1 Introduction
Discrete diffusion models (Austin et al., 2021; Campbell et al., 2022; Sun et al., 2023; Lou et al., 2024) and discrete flow-based models (Campbell et al., 2024; Gat et al., 2024; Shaul et al., 2025; Qin et al., 2025) have received significant attention for generating samples in discrete state spaces, providing effective alternatives to autoregressive (AR) models. Additionally, several works aim to study how to guide a pre-trained discrete model towards a desired distribution (Vignac et al., 2023; Schiff et al., 2025; Nisonoff et al., 2025). However, there are some challenges to developing guidance mechanisms in discrete cases. The main challenge is that discrete guidance is often associated with a transition probability or transition rate, which requires extra computation for all possible target positions after the transition. To reduce the number of forward passes through guidance models, existing methods treat the model as a continuous function and use the first-order approximation to compute guidance efficiently (Vignac et al., 2023; Schiff et al., 2025; Nisonoff et al., 2025). Unfortunately, this approximation might introduce non-negligible approximation errors. Moreover, existing literature only consider some particular cases, like class conditional guidance or energy-weighted guidance, which might not be general enough for various tasks.
We seek to develop a general guidance framework for discrete diffusion and flow-based models, which is illustrated in Fig. 1. Leveraging the Continuous-Time Markov Chain (CTMC) framework, given a pretrained model for sampling from a source distribution and the density ratio between the target distribution and source distribution, our method identifies the exact transition rate for sampling the target distribution. To the best of our knowledge, our guidance framework is the first exact discrete guidance in general form. Since our guidance formulation is expressed by the density ratio between the target distribution and source distribution, it also encompasses existing guidance methods as special cases. This is because the density ratio can be calculated by the energy function in energy-guided sampling, the ratio of the classifier in class-conditional generation, and preference alignment in reinforcement learning from human feedback (RLHF). In Table 1, we compare the formulation of the guidance and the number of function evaluations in each sampling step. Our framework only requires one forward pass in each sampling step for any initial distribution.
To compute the proposed guidance, we employ the Bregman divergence to train a network for estimating the conditional expectation. To further utilize potentially available samples from the target distribution, we further propose a regularization term.
@@
    We introduce a general guidance framework for discrete flow matching, achieving efficient sampling without approximation.
@@
    We propose to learn the guidance network by minimizing Bregman divergence. We further propose a regularization technique to utilize samples from the target distribution.
@@
    We verify the effectiveness of the proposed framework on energy-based sampling through simulations, preference alignment on text-to-image, and multimodal understanding benchmark.
Guidance Method Formula Exact Guidance # of function evaluations
@@
Rate-Based Nisonoff et al. (2025) Equation 9 $\mathcal{D}+1$ ; $\mathcal{D}\times({|\mathcal{S}|}-1)+1$
First-Order Approximated Nisonoff et al. (2025) Vignac et al. (2023); Schiff et al. (2025) Equation 13 2
  • Note: In Nisonoff et al. (2025), guidance requires $\mathcal{D}\times({|\mathcal{S}|}-1)+1$ function evaluations. When implemented using Algorithm 3, the number of function evaluations reduces to $\mathcal{D}+1$.
Table 1: Overall comparison with existing discrete guidance. Our posterior-based guidance provides an exact formulation of guidance for general target distributions with a sampling-efficient implementation. Existing methods are either limited to class-conditional generation or suffer from non-negligible approximation errors with multiple function evaluations in each sampling step. A more detailed discussion is provided in Section 3.
@@ 2 Discrete Flow Models
Discrete flow models (DFMs) aim to generate samples from a $\mathcal{D}$-dimensional discrete state-space with probability mass function (pmf) $q_{1}$, defined on the domain $\mathcal{S}^{\mathcal{D}}$, where $\mathcal{S}$ denotes the finite state space of each dimension. A natural example of this setting is text generation, where each coordinate corresponds to a token drawn from the vocabulary $\mathcal{S}$. To begin with, we first introduce some background on discrete flow models (Campbell et al., 2024; Gat et al., 2024; Shaul et al., 2025) under the framework of CTMC (Norris, 1998), which we describe in detail in Appendix C.
@@ 2.1 Transition Rate and Transition Probability
Thus, the transition rate $u_{t}^{q}$ is also the generator of the CTMC ${\{q_{t}\}}_{t\in[0,1]}$ (Holderrieth et al., 2025).
@@ 2.2 Modeling Transition Rate via Marginalization Trick
Given the transition rate $u_{t}^{q}(z,x)$ and current state $x$ at time $t$, we can generate a sample from $q_{t+h}$ using the equation 2, with a sufficiently small time step $h$. Therefore, we can learn a transition rate $u_{t}^{q}$ that can transport from an initial noise distribution $q_{0}$ to a target data distribution $q_{1}$. To obtain such a transition rate for sampling, a natural method is to learn the conditional expectation of the conditional transition rate $u^{q}_{t}(z,x|{\mathbf{x}}_{1})$ that generates the conditional probability path $q_{t|1}(\cdot|{\mathbf{x}}_{1})$ interpolating from noise to the datapoint ${\mathbf{x}}_{1}$. Then $u^{q}_{t}(z,x)=\mathbb{E}_{q_{1|t}({\mathbf{x}}_{1}|x)}{[u^{q}_{t}(z,x|{\mathbf{x}}_{1})]}$ can generate the target probability path $q_{t}$ (see Proposition 3.1 of Campbell et al., 2024), i.e., it satisfies the Kolmogorov forward equation 1. For completeness, we include the proof in Appendix D.1. We are free to define the conditional probability path $q_{t|1}(\cdot|{\mathbf{x}}_{1})$ and with the corresponding conditional transition rate as needed.
It is worth noting that learning a ${|\mathcal{S}|}^{\mathcal{D}}$-dimensional vector-valued function ${(u_{t}(z,x))}_{z\in\mathcal{\mathcal{S}}^{\mathcal{D}}}$ of current time $t$ and state $x$ is intractable when $\mathcal{D}$ is relatively large. To handle such a high-dimensional setting, a common approach is to construct a coordinate-wise conditional probability path and transition rate,
@@
which means that the elements of the vector ${\mathbf{x}}_{t}$ are independent conditional on ${\mathbf{x}}_{1}$. Here, $u_{t}^{q,d}(z^{d},x^{d}|x_{1}^{d})$ is the conditional transition rate that generates the conditional probability path $q_{t|1}^{d}$. A popular choice of probability path and the associated conditional transition rate used in the previous works (Campbell et al., 2024; Gat et al., 2024) is
@@
After defining the conditional path and rate, the marginal transition rate is given by
@@
where $q_{1|t}^{d}(x_{1}^{d}|x)=\sum_{x^{\backslash d}_{1}}q_{1|t}(x_{1}|x)$ is the posterior. Therefore, we only need to learn a $\mathcal{D}\times{|\mathcal{S}|}$-matrix-valued function $U_{t}^{q}(x)=(u_{t}^{q,d}(s,x))_{d,s}$. Then we can update each token in parallel during sampling.
@@
which is the cross-entropy between the true posterior $q_{1|t}$ and the estimated posterior $q^{\theta}_{1|t}$ (Campbell et al., 2024; Gat et al., 2024; Wang et al., 2025).
Additionally, there are some alternative approaches to training the transition rate, including directly maximizing the ELBO in terms of the unconditional transition rate (Shaul et al., 2025) and minimizing a Bregman divergence-type conditional generator matching loss (Holderrieth et al., 2025). It turns out that parameterization through the posterior gives us a unified framework to formulate the discrete guidance not only for discrete flow-based models, but also for other generative models. We include more detailed discussions in Section E.1. Moreover, details of the sampling implementation for equation 2 in Algorithm 1 are provided in the Appendix A.
@@ 3 Discrete Flow Guidance
In this section, we introduce our general guidance framework for modeling discrete data. It can be applied to both discrete flow models and discrete diffusion models (Campbell et al., 2022; Sun et al., 2023; Austin et al., 2021; Lou et al., 2024; Shi et al., 2024; Sahoo et al., 2024). More detailed discussions can be found in Appendix E.
@@ 3.1 Problem Formulation
Consider a CTMC with a probability path $p_{t}$ associated with a transition rate $u^{p}_{t}$. Suppose that the conditional transition rate and the conditional probability path are prespecified (see equation 4 for an example), and we are given a pre-trained posterior $p_{1|t}$ such that the unconditional transition rate $u^{p}_{t}(z,x)=\mathbb{E}_{{\mathbf{x}}_{1}\sim p_{1|t}({\mathbf{x}}_{1}|x)}{[u_{t}^{p}(z,x|{\mathbf{x}}_{1})]}$ can generate the probability path $p_{t}$. We will call the distribution with the pmf $p_{1}$ the source distribution. Our goal is to generate samples from the target distribution $q_{1}$ by rectifying the posterior or the unconditional transition rate corresponding to the probability path $p_{t}$.
@@ 3.2 Exact Discrete Guidance
Here, we only assume that the conditional probability paths are the same. Theorem 1 provides a general guidance framework for both discrete diffusion and flow models, because both of them can be modeled via parameterizing posterior densities. In discrete diffusion models, if we further assume that the transition rates of both forward noising processes are equal, then we can generalize and recover the predictor guidance in Nisonoff et al. (2025), which we state in the following.
@@
which recovers Equation 2 in Nisonoff et al. (2025).
@@
In general cases, the guidance is time-dependent according to Theorem 1 and Theorem 2. However, in the masked diffusion and flow models with $x_{1}$-independent scheduler, the guidance is time-independent. To see this, note that in those cases the posterior is time-independent (see Theorem 1 of Ou et al. (2025) and Proposition 6 of Gat et al. (2024) for details). Since the guidance is the integral of the product of the density ratio and the posterior, the time-independent posterior can lead to a time-independent guidance.
@@ 3.3 Training Objective
According to Theorem 1, given the density ratio $r(x)$, we can obtain the exact guidance term by learning the conditional expectation of the density ratio $h_{t}^{d}(z^{d},x)\overset{\triangle}{=}\mathbb{E}_{{\mathbf{x}}^{\backslash d}_{1}\sim p({\mathbf{x}}^{\backslash d}_{1}|{\mathbf{x}}^{d}_{1}=z^{d},{\mathbf{x}}_{t}=x)}{[r({\mathbf{x}}_{1})]}$. We parameterize the guidance $h_{t}$ using a neural network. A natural training objective for learning conditional expectation is the Bregman divergence (Banerjee et al., 2005), which is defined as
@@
where $F(\cdot)$ is a convex function. The most common choice of $F$ is $F(x)=\|{x}\|_{2}^{2}/2$, and then the Bregman divergence is the $\ell^{2}$-loss. Unfortunately, as mentioned in Lou et al. (2024), the $\ell^{2}$-loss does not work well for learning the conditional expectation of a density ratio due to the positive nature of the density ratio. In our work, we take $F(x)=\langle x,\log x\rangle$, and the training objective is given by the following.
@@
The loss $\mathcal{L}_{h,p}$ only requires the source distribution data. Additionally, we can utilize samples from the target distribution if they are available to achieve better performance. Similar to Ouyang et al. (2024), we introduce a regularization term allowing us to learn the guidance $h_{t}$ with sampling ${\mathbf{x}}_{1}\sim q_{1}$ during training. Notice that the minimizer of the objective in equation 5 is the underlying posterior $q_{1|t}$. Combining with Theorem 1, the exact guidance $h_{t}$ is the minimizer of the following objective.
@@ 3.4 Sampling
for each $d\in{[\mathcal{D}]}$, where $h_{t}^{d,\theta}(x)=\sum_{s\in\mathcal{S}}h_{t}^{d,\theta}(s,x)p_{1|t}^{d}(s,x)$. Then, we sample ${\mathbf{x}}^{d}_{t+h}$ by always-valid sampling procedure (see equation 14 in Appendix A for details). Specifically, at each step, we input the current state ${\mathbf{x}}_{t}$ into both the guidance network and pretrained posterior network, and then multiply the output to get the final transition probability in the target distribution $q^{d,\theta}_{1|t}(\cdot|{\mathbf{x}}_{t})$ to update the current state ${\mathbf{x}}_{t}$ like Algorithm 1. The overall sampling algorithm can be found in Algorithm 2 in Appendix A and the detailed design for the general rate-based method proposed in Theorem 2 can be found in Appendix F.
@@
The rate-based guidance involves multiple forward passes, which depend on the number of nonzero elements in the transition rate $u_{t}^{p}(z,{\mathbf{x}}_{t}),~z\neq{\mathbf{x}}_{t}$, given the current state ${\mathbf{x}}_{t}$. If one use the always-valid sampling algorithm with rate-based guidance Algorithm 3, the guidance entails $\mathcal{D}+1$ function calls in each sampling step, which is still computationally inefficient; see Fig. 2 (d). For predictor guidance, Nisonoff et al. (2025) proposed to use the first-order approximation to estimate the log-ratio $\log\frac{p(y|{\mathbf{x}}_{t}=z)}{p(y|{\mathbf{x}}_{t}=x)}$ for efficient sampling. To be specific, we give the general form based on our proposed rate-based guidance:
@@
3.5 Guidance for Preference Alignment
Given the general form of guidance for discrete flow matching, our framework is able to deal with energy-guided sampling (Zhang et al., 2025; Lu et al., 2023) and also preference alignment in reinforcement learning from human feedback (RLHF) (Rector-Brooks et al., 2025). In the following, we provide the formulation for RLHF. Let $\mathbf{c}$} denote the prompt with distribution $p_{\mathbf{c}}$, $\pi_{ref}({\mathbf{o}}_{1}|{\mathbf{c}})$ denote the reference policy associated with pretrained model $p^{d}_{1|t}({\mathbf{o}}_{1}^{d}|{\mathbf{c}},{\mathbf{o}}_{t})$. In the RL stage of RLHF (Ouyang et al., 2022), we consider to maximize
where $\tau$ is the temperature that controls the deviation from the reference policy. As mentioned in Rafailov et al. (2023), the above objective has the closed-form maximizer
where $\mathcal{Z}({\mathbf{c}})=\int\pi_{ref}({\mathbf{o}}_{1}|{\mathbf{c}})\exp{\Big(\frac{\mathcal{R}({\mathbf{c}},{\mathbf{o}}_{1})}{\tau}\Big)}\text{d}{\mathbf{o}}_{1}$ and $\mathcal{R}(\cdot,\cdot)$ is the reward function which can be rule-based or obtained by training a neural network on the a comparison dataset. Given the reward function, we aim to generate samples ${\mathbf{o}}_{1}\sim\pi^{*}({\mathbf{o}}_{1}|{\mathbf{c}})$ conditional on $\mathbf{c}$}. If we take $p_{1}(\cdot)=\pi_{ref}(\cdot|{\mathbf{c}})$ and $q_{1}(\cdot)=\pi^{*}(\cdot|{\mathbf{c}})$ in Theorem 1, we can obtain a rectified posterior for sampling from $\pi^{*}$, i.e., we would like to use $\mathbf{h}_{t}^{\theta}({\mathbf{o}}_{1},{\mathbf{o}}_{t},c)$ to approximate $\exp{\Big(\frac{\mathcal{R}({\mathbf{c}},{\mathbf{o}}_{1})}{\tau}\Big)}$. We leave the detailed training objective in equation G.2 in Appendix G.